# Can LLMs Handle WebShell Detection? Overcoming Detection Challenges with Behavioral Function-Aware Framework

### v3

@@ 2.1 Critical Function Filter
Our statistical analysis (detailed in Appendix B) shows that WebShell files use critical functions far more often than benign files. On average, WebShells contain 22.76 calls to critical functions, compared to only 0.74 in benign files, underscoring their behavioral complexity and distinctiveness. [^2] While this disparity in critical function usage is significant, it alone does not reliably distinguish WebShells from benign files, since legitimate scripts may also call these functions. Therefore, we use LLMs for deeper contextual analysis to improve detection accuracy.
@@ 2.4 LLM-Based Detection Framework
The input to the LLM consists of two main components: (a) a system directive that defines the model’s role as a cybersecurity expert, and (b) a user query that contains the extracted critical code segments and a selected ICL demonstration. To balance efficiency and performance, we limit the user query to one ICL demonstration, which reduces computational overhead while preserving sufficient context for reliable detection. Our prompt is described in detail in Appendix A.
@@ 3.1 Dataset Overview
We constructed a comprehensive dataset consisting of 21,665 benign PHP programs and 4,929 WebShells. The benign programs were obtained from established open-source PHP projects to ensure applicability to real-world scenarios. The WebShells were collected from public security repositories and augmented with synthetic obfuscation techniques to increase diversity.  [^3] We acknowledge the potential for data leakage due to the pre-training of LLMs. However, the autoregressive nature of this process does not explicitly capture WebShell classification, although it may affect other code generation tasks. To further minimize any risk of leakage, we restricted our selection of PHP programs to projects updated between October 2024 and 2025, thus reducing overlap with LLM training corpora. Using GPT-4’s tokenizer, we analyzed the token lengths of both sample types. The WebShell samples had a maximum token length of 1,386,438 and an average of 30,856.60 tokens, compared to a maximum of 305,670 tokens and an average of 2,242.89 tokens for benign programs. These results indicate that WebShells are typically significantly longer than benign samples. A detailed summary of the dataset composition can be found in the Table 3 in the Appendix C.
@@ 3.2 Experiment Setup
We randomly selected 60% of the dataset to create a fixed, known demonstration library for ICL. Using this subset, we computed normalized scores for different function categories based on the WBFP method, giving equal weight to coverage difference ($r_{c}$), frequency ratio ($r_{f}$), and usage ratio ($r_{u}$) to profile functions according to their behavioral importance in distinguishing WebShells from benign programs. These scores, detailed in the Table 4 in the Appendix C, guided the selection of ICL demonstrations from the known sample library.
@@ 4.2 The Effectiveness of Context-Aware Code Extraction (RQ2)
We evaluated the effectiveness of Context-Aware Code Extraction using two models: GPT-4, a large-scale model, and Qwen-2.5-3B, a smaller-scale model. Three configurations were compared: (1) predictions based on the full source code, (2) predictions using only extracted critical regions, and (3) a hybrid approach combining critical regions with truncated source code. Results are reported in Tables 5 and 6 in Appendix D.
@@ 4.3 The Effectiveness of WBFP for In-Context Learning (RQ3)
We evaluated the effectiveness of WBFP for ICL demonstration selection using Qwen-2.5-3B and GPT-4. This evaluation builds on the optimal hybrid strategy from RQ2. Four demonstration selection strategies were compared: Random Selection (Random), Source Code Semantic Similarity (SC-Sim), WBFP with Equal Weights (WBFP-Eq), and WBFP with Function-Level Weights (WBFP-Wt). Results are detailed in Tables 7 and 8 in Appendix D.
@@ 5 Related Work
6 Limitations and Future Work
Our research highlights several limitations that present important directions for future investigation in this emerging field.
6.1 Dataset Generalization
A fundamental challenge in WebShell detection is the scarcity of real-world data due to privacy concerns. Current research, including this study, mainly relies on public datasets, which may not fully represent the diversity and sophistication of novel threats. This reliance makes it difficult to rigorously evaluate a model’s ability to generalize to unseen attack vectors. This limits rigorous evaluation of a model’s ability to generalize against novel and heavily obfuscated threats. To address this, future work should focus on using LLMs to generate synthetic benchmarks. A large-scale, LLM-generated dataset would provide a privacy-preserving and diverse alternative for stress-testing model robustness against a wider range of attack vectors.
6.2 Model Robustness and System Adaptability
While our BFAD framework is effective, the underlying static analysis approach is vulnerable to sophisticated obfuscation and zero-day threats. Enhancing the core model and building adaptive deployment systems are critical next steps. To improve model robustness, future work could integrate dynamic analysis, examining runtime function call traces from sandboxed execution to unmask behavior hidden by obfuscation. Additionally, since our analysis revealed semantic understanding bottlenecks even in large models, domain-specific fine-tuning of code-centric LLMs on curated WebShell datasets is a crucial step to deepen their expertise and improve detection accuracy. For practical and resilient deployment, these enhanced models could be integrated into advanced architectures. A multi-stage ”fast and slow” system could use a smaller LLM for rapid, efficient filtering, while reserving a larger reasoning model for in-depth analysis of high-risk samples. To ensure long-term effectiveness against evolving threats, this pipeline could be governed by an autonomous agent that continuously identifies new attack patterns and automatically updates the detection system.
@@ 7 Conclusion
This paper presents a novel approach to WebShell detection using LLMs, addressing the unique challenges of applying these models to cybersecurity. We evaluated seven LLMs against traditional machine learning and deep learning methods using a dataset of 26.59K PHP scripts. Our analysis revealed that larger LLMs achieve near-perfect precision but lower recall, while smaller models exhibit high recall but poor precision, both of which underperform state-of-the-art methods such as GAT. To overcome these limitations, we proposed the BFAD framework, which integrates a critical function filter, context-aware code extraction, and WBFP. BFAD significantly improved LLM performance, with GPT-4 achieving an F1 of 99.35%, surpassing traditional benchmarks, and smaller models such as Qwen-2.5-0.5B improving by 51.23 percentage points to an F1 of 82.67%. As the first systematic exploration of LLMs for WebShell detection, this work provides actionable solutions to their contextual and behavioral challenges, paving the way for future advances in code security.
Appendix A Prompt Details
@@
Appendix B Critical Function Details
@@
Appendix C Dataset Details

### v4

@@ 2.1 Critical Function Filter
Our statistical analysis (detailed in Appendix C) shows that WebShell files use critical functions far more often than benign files. On average, WebShells contain 22.76 calls to critical functions, compared to only 0.74 in benign files, underscoring their behavioral complexity and distinctiveness. [^2] While this disparity in critical function usage is significant, it alone does not reliably distinguish WebShells from benign files, since legitimate scripts may also call these functions. Therefore, we use LLMs for deeper contextual analysis to improve detection accuracy.
@@ 2.4 LLM-Based Detection Framework
The input to the LLM consists of two main components: (a) a system directive that defines the model’s role as a cybersecurity expert, and (b) a user query that contains the extracted critical code segments and a selected ICL demonstration. To balance efficiency and performance, we limit the user query to one ICL demonstration, which reduces computational overhead while preserving sufficient context for reliable detection. Our prompt is described in detail in Appendix B.
@@ 3.1 Dataset Overview
We constructed a comprehensive dataset consisting of 21,665 benign PHP programs and 4,929 WebShells. The benign programs were obtained from established open-source PHP projects to ensure applicability to real-world scenarios. The WebShells were collected from public security repositories and augmented with synthetic obfuscation techniques to increase diversity.  [^3] We acknowledge the potential for data leakage due to the pre-training of LLMs. However, the autoregressive nature of this process does not explicitly capture WebShell classification, although it may affect other code generation tasks. To further minimize any risk of leakage, we restricted our selection of PHP programs to projects updated between October 2024 and 2025, thus reducing overlap with LLM training corpora. Using GPT-4’s tokenizer, we analyzed the token lengths of both sample types. The WebShell samples had a maximum token length of 1,386,438 and an average of 30,856.60 tokens, compared to a maximum of 305,670 tokens and an average of 2,242.89 tokens for benign programs. These results indicate that WebShells are typically significantly longer than benign samples. A detailed summary of the dataset composition can be found in the Table 3 in the Appendix D.
@@ 3.2 Experiment Setup
We randomly selected 60% of the dataset to create a fixed, known demonstration library for ICL. Using this subset, we computed normalized scores for different function categories based on the WBFP method, giving equal weight to coverage difference ($r_{c}$), frequency ratio ($r_{f}$), and usage ratio ($r_{u}$) to profile functions according to their behavioral importance in distinguishing WebShells from benign programs. These scores, detailed in the Table 4 in the Appendix D, guided the selection of ICL demonstrations from the known sample library.
@@ 4.2 The Effectiveness of Context-Aware Code Extraction (RQ2)
We evaluated the effectiveness of Context-Aware Code Extraction using two models: GPT-4, a large-scale model, and Qwen-2.5-3B, a smaller-scale model. Three configurations were compared: (1) predictions based on the full source code, (2) predictions using only extracted critical regions, and (3) a hybrid approach combining critical regions with truncated source code. Results are reported in Tables 5 and 6 in Appendix E.
@@ 4.3 The Effectiveness of WBFP for In-Context Learning (RQ3)
We evaluated the effectiveness of WBFP for ICL demonstration selection using Qwen-2.5-3B and GPT-4. This evaluation builds on the optimal hybrid strategy from RQ2. Four demonstration selection strategies were compared: Random Selection (Random), Source Code Semantic Similarity (SC-Sim), WBFP with Equal Weights (WBFP-Eq), and WBFP with Function-Level Weights (WBFP-Wt). Results are detailed in Tables 7 and 8 in Appendix E.
@@ 6 Conclusion
This paper systematically demonstrates that baseline LLMs are suboptimal for WebShell detection, exhibiting a precision-recall trade-off that underperforms state-of-the-art methods. We propose the Behavioral Function-Aware Detection (BFAD) framework, which overcomes these limitations by focusing analysis on critical code segments and optimizing in-context learning. BFAD enables our LLM-based approach to surpass traditional benchmarks.
Appendix A Limitations and Future Work
Our research also highlights several limitations that present important directions for future investigation in this emerging field.
A.1 Dataset Generalization
A fundamental challenge in WebShell research is the scarcity of diverse, real-world data. Our reliance on public datasets, like that of prior work, limits the evaluation of model generalization against novel, sophisticated threats. To address this, future work should focus on using LLMs to generate large-scale synthetic benchmarks. A large-scale, LLM-generated dataset would provide a privacy-preserving and diverse alternative for stress-testing model robustness against a wider range of attack vectors.
A.2 Model and System Robustness
Our framework’s static analysis approach remains vulnerable to advanced obfuscation and zero-day attacks. Future work should therefore focus on enhancing model robustness and system adaptability. To counter evasion, models could integrate dynamic analysis by examining runtime function call traces from sandboxed execution. To deepen semantic understanding, domain-specific fine-tuning of code-centric LLMs on curated WebShell datasets is a crucial next step. For practical deployment, these enhanced models could be integrated into a multi-stage ”fast and slow” system, using a smaller model for rapid filtering and a larger one for in-depth analysis of high-risk samples. To ensure long-term resilience, this pipeline could be governed by an autonomous agent that continuously identifies new attack patterns and automatically updates the detection models.
Appendix B Prompt Details
@@
Appendix C Critical Function Details
@@
Appendix D Dataset Details